# Multi-Modal Diffusion for Hand-Object Grasp Generation

### v1

@@ 3.3.1 Latent Codes for Different Modalities
For each modality aforementioned, we have an MLP-based encoding network to convert the raw representation into latent codes:
@@
with the dimensions $\mathbf{z}_{0}^{H}\in\mathbb{R}^{1\times 512}$ amd $\mathbf{z}_{0}^{O}\in\mathbb{R}^{4\times 512}$ They are concatenated to be the final latent code for the whole hand-object interaction configuration
@@
This design allows us to jointly process different modalities without unnecessary entanglement. This allows us to train the model with partial modality supervision without corrupting the parameters for other modalities.
@@ 3.3.2 Asynchronous Denoising Schedulers
We propose to generate either both hand and object or only the hand given an object shape. To realize this, we need a certain degree of entanglement between the two modalities as we need them to cooperate to form a valid grasp. However, we still desire certain isolation so that we could supervise the model training with partially annotated data, for example only the 3D object shapes. In the usual fashion of organizing multi-modal latent codes in latent diffusion models, the latent codes of different modalities always have the same level of noise corruption as they share a single noise scheduler. However, we desire the two modalities can be denoised with respect to an arbitrary level of noise in the other modality. Therefore, we propose to use asynchronous denoising schedulers for this purpose during training. Instead of using a single noise scheduler, we have two schedulers $t_{H}$ and $t_{O}$ to control the noise patterns in the hand latent code and the object latent code respectively. Following Equation 3 we derive the corrupted latent codes as
$\mathbf{z}^{H}_{t_{H}}=\Pi_{t=1}^{t_{H}}q(\mathbf{z}^{H}_{t}|\mathbf{z}^{H}_{t-1}),\quad\mathbf{z}^{O}_{t_{O}}=\Pi_{t=1}^{t_{O}}q(\mathbf{z}^{O}_{t}|\mathbf{z}^{O}_{t-1}),$ (9)
with the same time step range $t_{O},t_{H}\in[1,T]$. This design enables the diffusion to deals with the noise in each modality separately. During training, it allows using object-only data to supervise the object part only. During sampling, it allows the model to generate only the hand given an object shape as the condition. Similar to Bao et al. [3], we will learn the joint distribution for hand-object pairs and the marginal distribution for the hand at the same time in an end-to-end fashion by manipulating the time schedulers.
@@ 3.3.3 Training and Sampling
Now we introduce the training and sampling paradigm of MGD. There are two main goals in designing it. First, we would like to use the training with heterogeneous annotations so that we can go beyond the limited object shapes in HOI datasets. Second, we want a single model capable of object-conditioned hand grasp generation and unconditional hand-object joint generation.
Training. When training on the canonical HOI data samples, with the asynchronous noise schedulers, we can derive the assembly of two corrupted latent codes as
@@
Due to the element-wise property of MSE loss used for diffusion model training, we can calculate and back-propagate the gradient to a certain part of the latent codes independently. We leverage this property to train the unconditional and conditional generation at the same time. First, we learn the joint distribution by the unconditional generation loss
$L_{\text{uncond.}}=\mathbb{E}_{z_{0}\sim\mathcal{D}_{HO},\{t_{H},t_{O}\}\sim[1,T]}\Bigr{[}||z_{0}-G_{\Phi}(z_{t_{H},t_{O}},t_{H},t_{O})||^{2}_{2}\Bigr{]},$ (11)
where $\mathcal{D}_{HO}$ is the distribution of the HOI datasets [26, 33].
Also, we train the model to learn the object-conditioned grasp generation. In this situation, the object latent code is given and fixed as $\mathbf{z}_{0}^{O}$ and we set $t_{O}=0$ indicating that the object part is noise-free and serves as a condition to derive the marginal distribution now. We now train with the loss
$L_{\text{cond.}}=\mathbb{E}_{z_{0}\sim\mathcal{D}_{HO},t_{H}\sim[1,T]}\Bigr{[}||z_{0}-G_{\Phi}(z_{t_{H}},t_{H})||^{2}_{2}\Bigr{]},$ (12)
with
$t_{O}=0,\quad z_{t_{H}}=z_{t_{H}}^{H}\oplus\mathbf{z}_{0}^{O}.$ (13)
The conditional and unconditional generation losses require uncorrupted and corrupted object latent codes respectively. In practice, we combine these two training objectives in a 1:1 ratio for a single draw of the training data batch. This training strategy is similar to the classifier-free guidance [10] for diffusion model training. However, the optional condition (object) exists in the input and output latent instead of in an independent condition vector.
Finally, we could leverage the large-scale 3D object shape datasets to train the object distribution only. This is done by selecting object samples from a wider source of datasets as we do not need available hand grasps over them anymore. The loss turns to
$L_{\text{obj.}}=\mathbb{E}_{z_{0}^{O}\sim\mathcal{D}_{O}\cup\mathcal{D}_{HO},t_{O}\sim[1,T]}\Bigr{[}||z_{0}-G_{\Phi}(z_{t_{O}},t_{O})||^{2}_{2}\Bigr{]},$ (14)
where $\mathcal{D}_{O}$ is object-only datasets. The latent codes only take the object part as the variable:
@@
Obviously, the gradient only influences the object-related parts while keeping the hand part as-is. During training, the object encoder and decoder networks are pre-trained on the ShapeNet [4] and frozen. All other encoder, decoder, and denoiser networks are trained end to end.
Similar to the proximity sensor features [27] and the Grasping Filed [13], we use a distance field to help capture the relation between the hand and the object. Since we eliminate the requirement of object templates to accommodate more general datasets, we can not define the field using object keypoints [7] or templated vertices. Instead, we define the distance field as the vectors between each hand joint and the 10 closest points in the object points. Therefore, we have the distance field as $\mathbf{y}_{f}\in\mathbb{R}^{16\times 30}$. When the data sample is drawn from HOI datasets, we would also supervise the distance field calculated from the recovered hand gesture and object shape:
$L_{\text{distance}}=||\mathbf{y}_{f}-\hat{\mathbf{y}}_{f}||_{2}^{2},$ (16)
@@
$L_{\text{hand\_xyz}}=||\text{LBS}(\mathbf{y}_{H})-\text{LBS}(\hat{\mathbf{y}}_{H})||_{2}^{2},$ (17)
where $\text{LBS}(\cdot)$ is the linear blend skinning process by MANO parametric model to derive the joint kinematic positions. With all the losses introduced, we could train the multi-modal diffusion model using data with heterogeneous annotations and for unconditional and conditional generation at the same time. The overall loss is
$L=\mathbbm{1}_{z_{0}\sim D_{HO}}(L_{\text{uncond.}}+L_{\text{cond.}}+L_{\text{distance}}+L_{\text{hand\_xyz}})+L_{\text{obj.}},$ (18)
where $\mathbbm{1}_{z_{0}\sim D_{HO}}$ indicates whether the data sample is drawn from a HOI dataset.
Sampling. We follow the canonical progressive denoising process for diffusion models in the sampling stage. To sample for joint (unconditional) HOI generation, we synchronize the schedulers $t_{H}=t_{O}=t$. Similar to Tevet et al. [28], at each step $t$, we predict a clean sample by $\hat{z}_{0}=G_{\Phi}(z_{t},t_{H},t_{O})$ from the corrupted code $z_{t}$ and then noise it back to $z_{t-1}$. During sampling for object-conditioned grasp generation, we keep $t_{O}=0$ and $z_{t_{H}}=z_{t_{H}}^{H}\oplus z_{0}^{O}$. Then the denoiser predicts $\hat{z}_{0}=G_{\Phi}(z_{t_{H}},t_{H})$ and then noise it back to $z_{t-1}$. In either the unconditional or the conditional generation, we repeat the process above along $t=T\longrightarrow 1$ until the final $\hat{z}_{0}$ is achieved after $T$ iterations.
@@ 4 Experiments
Table 1: Quantitative evaluation of grasp generation. The top four entries are evaluated on objects from the OakInk-test set. The bottom two entries are evaluated on the same set of objects generated from MGDin the unconditional mode, which are out of the domain from the objects seen during training.
Objects Methods FID $\downarrow$ Pene. Dep. $\downarrow$ Intsec. Vol. $\downarrow$ Sim. Disp. Mean $\downarrow$ User Scor.$\uparrow$
OakInk-test GrabNet [26] 26.96 0.70 11.93 3.14 3.80
GrabNet-Refine [26, 33] 25.26 0.63 4.44 2.78 4.00
MGD 20.73 0.32 3.87 2.52 4.87
@@
MGD(uncond.) - 0.51 6.71 7.13 3.53
Table 2: The ablation study about the impact of training data on the generation quality over unseen objects from ARCTIC [7]. Adding more training data consistently boosts the generation quality on unseen objects.
OakInk-Shape GRAB Object-only Data FID $\downarrow$ Pene. Dep. $\downarrow$ Intsec. Vol. $\downarrow$ Sim. Disp. Mean $\downarrow$ User Scor.$\uparrow$
@@
OakInk-Shape GRAB Object-only Data $\downarrow$ Pene. Dep. $\downarrow$ Intsec. Vol. $\downarrow$ Sim. Disp. Mean $\downarrow$ User Scor.$\uparrow$
@@
Table 4: Ablation about asynchronous denoising schedulers. We evaluate on the generated object shapes by MGD.
@@ 4.1 Setups
Datasets. The hand-object interaction (HOI) datasets usually face limitations of the diversity of object shapes. We combine the data from multiple resources to train the model. GRAB [26] contains human full-body poses together with 3D objects. We extract the hands from the full-body annotations. For OakInk-Shape [33], we use the official training split for training. We also use the contact-adapted synthetic grasp on the OakInk-Shape set for training. Besides the hand-object interaction data, we also leverage the rich resources of 3D object data to help train the object part in our model. The object data is also used to train the LION code prior distribution as the original distribution (pre-trained from ShapeNet [4]) can produce object shapes that are not proper for single-hand grasp, for example, sofa, chair, and bookcase. We combine the objects from GRAB, OakInk (both the objects from OakInk-Image with grasp annotation and the objects from OakInk-shape without grasps), Affordpose [35] and DexGraspNet [30] as the object data resource. We also leverage the DeepSDF [18] as trained in Ink-base [33] to provide synthetic object data and include them in OakInk-Shape. To evaluate the generation quality on unseen objects, we leave the ARCTIC [7] dataset away from training and only used for evaluation purposes. Please refer to the supplemental for the statistics of each dataset.
@@
Data Pre-Processing. We follow the convention in OakInk-Shape to transform the hands into the object-centric frame. For data from GRAB [26], we extract the MANO parameters from its original SMPL-X [19] annotations. Then, we filter out the frames where the right-hand mesh has no contact with the object in training. For the object representation, we randomly sample 2048 points from mesh vertices to represent each object. For the objects in the OakInk-Shape dataset, we sample the point clouds from the object mesh surfaces uniformly instead of sampling from the vertices because the annotated mesh vertices are too sparse on smooth surfaces.
@@
Implementation. The DeepSDF in Ink [33] is trained per category to derive better details and fidelity when interpolating object shapes. We use the DeepSDF to provide synthetic object shapes. For the tokenizer network (transforming modality features to latent codes), we always use 2-layer MLP networks with a hidden dimension of 1024 and an output dimension of 512. For the denoiser, we follow MDM [28] to use a transformer encoder-only backbone-based diffusion network. After decoding the object latent code by the LION decoder, we derive the point cloud set. For visualization purposes, we conduct surface reconstruction from the generated object point cloud. Though advanced surface reconstruction techniques such as some learned solver [20] can provide more details, this part is not our focus and we wish to keep consistency for out-of-domain objects generated. Therefore, we choose the naive Alpha Shape algorithm [6].
@@
Evaluation and Metrics. Unlike the reconstruction tasks where the ground truth is available, it is difficult to do the quantitative evaluation for generation models. We keep the OakInk-Shape test set for quantitative evaluations and follow its metrics set. Referring to the evaluation of image generation quality, we first measure the quality of generated grasp by FID (Frechet Inception Distance) [9] between the images rendered from the ground truth grasps and the generated grasps. We can not directly calculate the FID score on generated 3D parameters as there is no commonly used encoder for this purpose and implementing it ourselves can cause many ambiguities. We also measure the model performance by directly checking the 3D asset quality. We follow previous studies [33, 32] to use three metrics: (1) the penetration depth (Pene. Depth) between hand and object mesh, (2) the solid intersection volume (Intsec. Vol.) between them and (3) the mean displacement in simulation following [8]. Finally, we perform a user study to measure the plausibility and the truthfulness of the generated grasps by scoring (1-5). We invite 15 participants to rate the quality of batches of the mixture of 10 rendering of ground truth grasps and 10 generated grasps. The score 1 indicates “totally fake and implausible” and 5 indicates “plausible enough to be real”. Each participant evaluates by averaging the scores over 10 randomly selected batches.

### v2

@@ 3.3.1 Latent Codes for Different Modalities
For each modality, we have an MLP-based encoding network to convert the raw representation into latent codes:
@@
with the dimensions $\mathbf{z}_{0}^{H}\in\mathbb{R}^{1\times 512}$ amd $\mathbf{z}_{0}^{O}\in\mathbb{R}^{4\times 512}$ . They are concatenated to be the final latent code for the whole hand-object interaction configuration
@@
This design allows us to jointly process different modalities without unnecessary entanglement. So we can train the model with partial modality supervision without corrupting the parameters for other modalities.
@@ 3.3.2 Asynchronous Denoising Schedulers
We propose to generate either both hand and object or only the hand given an object shape. To realize this, we need a certain degree of entanglement between the two modalities as we need them to cooperate to form a valid grasp. However, we still desire certain disentanglement so that we could supervise the training with partially annotated data, e.g., 3D object shape data. In the usual fashion of organizing multi-modal latent codes in latent diffusion models, the latent codes of different modalities always have the same scale of noise corruption as they share the noise scheduler. In contrast, we desire that the two modalities can be denoised with respect to an arbitrary level of noise in the other modality. Therefore, we propose to use asynchronous denoising schedulers for this purpose during training. Instead of using a single noise scheduler, we have two schedulers $t_{H}$ and $t_{O}$ to control the noise patterns in the hand latent code and the object latent code respectively. Following Equation˜3 we derive the corrupted latent codes as
$\mathbf{z}^{H}_{t_{H}}=\Pi_{t=1}^{t_{H}}q(\mathbf{z}^{H}_{t}|\mathbf{z}^{H}_{t-1},\mathbf{z}^{O}_{t-1}),\quad\mathbf{z}^{O}_{t_{O}}=\Pi_{t=1}^{t_{O}}q(\mathbf{z}^{O}_{t}|\mathbf{z}^{O}_{t-1}),$ (9)
with the same time step range $t_{O},t_{H}\in[1,T]$. This design allows the diffusion to deal with the noise in each modality separately. During training, it allows using object-only data to supervise the object part only. During sampling, it allows the model to generate only the hand given an object shape as the condition. Similar to Bao et al. (2023), we will learn the joint distribution for hand-object pairs and the marginal distribution for the hand at the same time in an end-to-end fashion by manipulating the time schedulers.
@@ 3.3.3 Training and Sampling
Training. When training on HOI data samples, with the asynchronous noise schedulers, we can derive the assembly of two corrupted latent codes as
@@
Due to the element-wise property of MSE loss, we can calculate and back-propagate the gradient to a certain part of the latent codes independently. We leverage this property to train the unconditional and conditional generation at the same time. To learn the joint distribution, we apply the unconditional generation loss
$L_{\text{uncond.}}=\mathbb{E}_{z_{0}\sim\mathcal{D}_{HOI},\{t_{H},t_{O}\}\sim[1,T]}\Bigr[||z_{0}-G_{\Phi}(z_{t_{H},t_{O}},t_{H},t_{O})||^{2}_{2}\Bigr],$ (11)
where $\mathcal{D}_{HOI}$ is the distribution of the HOI datasets (Taheri et al., 2020; Yang et al., 2022). We also train the model to learn the object-conditioned grasp generation. Given an object latent code $\mathbf{z}_{0}^{O}$, we set $t_{O}=0$ indicating that the object part is noise-free and serves as a condition to derive the marginal distribution. The object-conditioned generation loss is thus
$L_{\text{cond.}}=\mathbb{E}_{z_{0}\sim\mathcal{D}_{HOI},t_{H}\sim[1,T]}\Bigr[||z_{0}-G_{\Phi}(z_{t_{H}},t_{H})||^{2}_{2}\Bigr],$ (12)
with $z_{t_{H}}=z_{t_{H}}^{H}\oplus\mathbf{z}_{0}^{O}$. The conditional and unconditional generation losses require uncorrupted and corrupted object latent codes respectively. In practice, we combine these two training objectives in a 1:1 ratio for a single draw of training data batch. This training strategy is similar to the classifier-free guidance (Ho and Salimans, 2022) for diffusion model training. However, the optional condition (object) exists in the input and output instead of in an independent condition vector.
Finally, we could leverage the large-scale 3D object shape datasets to train the object distribution only. This learning goal is designed to improve the object latent code by exposing it to a broader distribution of object shapes. We select object samples from both HOI datasets $\mathcal{D}_{HOI}$ and object-only datasets $\mathcal{D}_{O}$. The loss turns to
$L_{\text{obj.}}=\mathbb{E}_{z_{0}^{O}\sim\mathcal{D}_{O}\cup\mathcal{D}_{HOI},t_{O}\sim[1,T]}\Bigr[||z_{0}-G_{\Phi}(z_{t_{O}},t_{O})||^{2}_{2}\Bigr],$ (13)
where we apply no supervision on the hand branch. While a hand latent code is present as a part of the joint latent code, it is randomly initialized and remains untrained. The latent codes only take the object part as the variable:
@@
Obviously, the gradient only influences the object-related parts while keeping the hand part as-is. During training, the object encoder and decoder networks are pre-trained on the ShapeNet (Chang et al., 2015) and frozen. All other encoder, decoder, and denoiser networks are trained end to end.
Similar to the proximity sensor features (Taheri et al., 2024) and the Grasping Filed (Karunratanakul et al., 2020), we use a distance field to help capture the relation between the hand and the object. Since we eliminate the requirement of object templates to accommodate more general datasets, we can not define the field using object keypoints (Fan et al., 2023) or templated vertices. Instead, we define the distance field as the vectors between each hand joint and the 10 closest points in the object points. Therefore, we have the distance field as $\mathbf{y}_{f}\in\mathbb{R}^{16\times 30}$. When the data sample is drawn from HOI datasets, we would also supervise the distance field calculated from the recovered hand gesture and object shape:
[…]
@@
[…]
where $\text{LBS}(\cdot)$ is the linear blend skinning process by MANO parametric model to derive the joint kinematic positions. With all the losses introduced, we could train the diffusion model using data with heterogeneous annotations and for unconditional and conditional generation at the same time. The overall loss is
$L=\mathbbm{1}_{z_{0}\sim D_{HOI}}(L_{\text{uncond.}}+L_{\text{cond.}}+L_{\text{distance}}+L_{\text{hand\_xyz}})+L_{\text{obj.}}.$ (17)
Sampling. We follow the canonical progressive denoising process for diffusion models in the sampling stage. To sample for joint (unconditional) HOI generation, we synchronize the schedulers $t_{H}=t_{O}=t$. Similar to Tevet et al. (2022), at each step $t$, we predict a clean sample by $\hat{z}_{0}=G_{\Phi}(z_{t},t_{H},t_{O})$ from the corrupted code $z_{t}$ and then noise it back to $z_{t-1}$. During sampling for object-conditioned grasp generation, we keep $t_{O}=0$ and $z_{t_{H}}=z_{t_{H}}^{H}\oplus z_{0}^{O}$. Then the denoiser predicts $\hat{z}_{0}=G_{\Phi}(z_{t_{H}},t_{H})$ and then noise it back to $z_{t-1}$. In either the unconditional or the conditional generation, we repeat the process above along $t=T\longrightarrow 1$ until the final $\hat{z}_{0}$ is achieved after $T$ iterations.
@@ 4 Experiments
Table 1: Quantitative evaluation of grasp generation. All models are trained using GRAB and OakInk-shape training set. We evaluate the methods on the objects from OakInk-test set, objects generated by JHOD, and objects from the ARCTIC dataset.
[…]
OakInk-test GrabNet 26.96 0.70 11.93 3.14 3.80
GrabNet-Refine 25.26 0.63 4.44 2.78 4.00
JHOD 20.73 0.32 3.87 2.52 4.87
ARCTIC GrabNet-Refine 37.23 1.22 14.20 12.17 1.33
JHOD 23.92 0.42 4.55 2.91 4.13
@@
JHOD(uncond.) - 0.51 6.71 7.13 3.53
Table 2: The ablation study about the impact of training data on the generation quality over unseen objects from ARCTIC (Fan et al., 2023). Adding more training data consistently boosts the generation quality on unseen objects.
[…]
@@
[…]
@@
Table 4: Ablation about asynchronous denoising schedulers. We evaluate on the generated object shapes by JHOD.
@@ 4.1 Setups
Datasets. We combine the data from multiple resources to train the model. GRAB (Taheri et al., 2020) contains human full-body poses together with 3D objects. We extract the hands from the full-body annotations. For OakInk (Yang et al., 2022), we use the official training split for training. We also use the contact-adapted synthetic grasp from the OakInk-Shape dataset for training. Besides the hand-object interaction data, we also leverage the rich resources of 3D object data to help train the object part in our model. The object data is also used to fine-tune the LION prior distribution. Fine-tuning prevents improper object shapes from the original distribution (pre-trained from ShapeNet (Chang et al., 2015)) such as sofas, chairs, and bookcases. We combine the objects from GRAB, OakInk (both the objects from OakInk-Image with grasp annotation and the objects from OakInk-shape without grasps), Affordpose (Ye et al., 2023b) and DexGraspNet (Wang et al., 2023) as the object data resource. We also leverage the DeepSDF (Park et al., 2019) as trained in Ink-base (Yang et al., 2022) to provide synthetic object data and include them in OakInk-Shape.
@@
Data Pre-Processing. We follow the convention in OakInk-Shape to transform the hands into the object-centric frame. For data from GRAB (Taheri et al., 2020), we extract the MANO parameters from its original SMPL-X (Pavlakos et al., 2019) annotations. Then, we filter out the frames where the right-hand mesh has no contact with the object in training. For the object representation, we randomly sample 2048 points from mesh vertices to represent each object every time we draw the object in both training and inference to avoid overfitting a fixed set of point clouds. For the objects in the OakInk-Shape dataset, we sample the point clouds from the object mesh surfaces uniformly instead of sampling from the vertices because the annotated mesh vertices are too sparse on smooth surfaces.
@@
Implementation. The DeepSDF in Ink (Yang et al., 2022) is trained per category to derive better details and fidelity when interpolating object shapes. We use the DeepSDF to provide synthetic object shapes. For the encoder network to transform modality features to latent codes, we always use 2-layer MLP networks with a hidden dimension of 1024 and an output dimension of 512. For the denoiser, we follow MDM (Tevet et al., 2022) to use a transformer encoder-only backbone-based diffusion network. After decoding the object latent code by the LION decoder, we derive the point cloud set. For visualization purposes, we conduct surface reconstruction from the generated object point cloud. Though advanced surface reconstruction techniques such as some learned solver (Peng et al., 2021) can provide more details, this part is not our focus and we wish to keep consistency for out-of-domain objects generated. Therefore, we choose the classic Alpha Shape algorithm (Edelsbrunner et al., 1983). Without category labels or other constraints, it is challenging to generate object shapes with highly carved details and it is not our focus in this paper.
Baseline Methods There is no commonly adopted benchmark in the area of grasp generation and some similar methods can follow different evaluation protocols. For example, the provided evaluation protocol of G-HOP (Ye et al., 2024) is for grasp reconstruction instead of generation and text description or object label is necessary for its generation mode which is not required in our proposed method. On the other hand, another line of works, such as UGG (Lu et al., 2023) and DexDiffuer (Weng et al., 2024) focus on grasp generation but within the domain of robotic dexterous hand thus there is no trivial way to compare with their in the same setting. In this work, we focus on hand grasp generation with only object shape as condition or without any condition. We select the widely adopted method GrabNet (Taheri et al., 2020) as the baseline method to compare with in this section. We also use its refined version GrabNet-Refine in some certain comparisons.
@@
Evaluation and Metrics. Unlike the reconstruction tasks where the ground truth is available, it is difficult to do the quantitative evaluation for generation models. We hold the objects from the OakInk-Shape test set and ARCTIC (Fan et al., 2023) dataset for quantitative evaluations. We measure the quality of generated grasp by FID (Frechet Inception Distance) (Heusel et al., 2017) between the images rendered from the ground truth grasps and the generated grasps. We note that we can not directly calculate the FID score on generated 3D parameters as there is no commonly used encoder for this purpose and implementing it ourselves can cause many ambiguities. Therefore, we follow the practice Chan et al. (2022); Gao et al. (2022) of computing metrics on rendered images. Also, as the variance of camera views can significantly impact the image distribution, we render three views of each GT or generated grasp. We render three views because penetration or implausible contact may only be visible from certain angles. This multi-view rendering can provide more comprehensive qualitative evaluation but also reduces the potential bias introduced by the setup of single camera view. We also measure the model performance by directly checking the 3D asset quality. We follow previous studies (Yang et al., 2022; 2021) to use three metrics: (1) the penetration depth (Pene. Depth) between hand and object mesh, (2) the solid intersection volume (Intsec. Vol.) between them and (3) the mean displacement in simulation following (Hasson et al., 2019). Finally, we perform a user study to measure the plausibility and the truthfulness of the generated grasps by scoring (1-5). We invite 15 participants to rate the quality of batches of the mixture of 10 rendering of ground truth grasps and 10 generated grasps. The score 1 indicates “totally fake and implausible” and 5 indicates “plausible enough to be real”. Each participant evaluates by averaging the scores over 10 randomly selected batches.